OpenStudy (anonymous):

plz help me i really need this credit. thanx in advance:) Find the missing term. 7^-147 × 7^98 = 7^18 × 7^-38 ×????

OpenStudy (welshfella):

add the exponents on left side we have - 147 + 98 = -49 7^-49 now the right side should also be 7 ^-49 the exponents should add up to -49

OpenStudy (anonymous):

Sorry im not quite sure what to do with the -49 but the answer choices are: A. 7^(-20) B. 7^22 C. 7^(-29) D. 7^58

OpenStudy (anonymous):

I think the answer is 7^(-20)

OpenStudy (anonymous):

thank you welshfella i get it now

OpenStudy (welshfella):

18 + -38 + x = -49 find x and you have the answer

OpenStudy (welshfella):

x = -49 - 18 + 38 = ?

OpenStudy (welshfella):

its not -20

OpenStudy (anonymous):

I know im goiing to try to solve the equation and see what i get

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

so its -29 then?

OpenStudy (welshfella):

yep

OpenStudy (welshfella):

recall when you multiply expeontial terms you ADD the exponents eg 5^2 * 5^3 = 5^5

OpenStudy (welshfella):

answer is 7^-29

OpenStudy (anonymous):

yes i'll remember that. thank you so much

OpenStudy (welshfella):

yw